Shyam and Haru's present total age is 50 yrs. 20 yrs before the age of Haru is twice than shyam. What is the present age of Haru and Shyam?

Answer by stanbon(75887)

Shyam and Haru's present total age is 50 yrs. 20 yrs before the age of Haru is twice than shyam. What is the present age of Haru and Shyam?
-----
Equations:
s + h = 50
h-20 = 2(s-20)
----
Rearrange:
s + h = 50
2s- h = 20
----
3s = 70
s = 23 yrs 4 months
------------------------------
h = 50 - 23.33 = 26 yrs 8 months
=================================
Cheers,
Stan H.
============